OpenStudy (anonymous):

how do you simplify 2x-2y+5z-2x-y+3z

OpenStudy (cwrw238):

you can only simplify like terms so bring like terms together that is 2x - 2x -2y - y +5z + 3z so for example 5z + 3z = 8z now do same for terms in x and y

OpenStudy (anonymous):

2x - 2y + 5z - 2x - y + 3z - 2y + 5z - y + 3z (2x-2x=0) - 2y + 8z - y (3z+5z=8z) - 3y + 8z (-2y-y) 8z - 3y.

OpenStudy (anonymous):

2x-2y+5z-2x-y+3z collect like terms 5z+3z-2y-y+2x-2x 2x-2x will cancel out therefore 5z+3z-2y-y=8z-3y .